# Multi-modal AI for comprehensive breast cancer prognostication

### v1

@@ 4.3.1 Creating model ensembles
Ensembling the clinical models happened in two steps. First, the output risk scores from all the clinical models were scaled approximately to the interval $[0,1]$. That is, for each model $m_{k}$, we found the maximum over the entire validation set, $\mathrm{max}^{k}=\mathrm{max}_{i}\{\hat{y}^{k}_{i}\}$, and the minimum over the entire validation set, $\mathrm{min}^{k}=\mathrm{min}_{i}\{\hat{y}^{k}_{i}\}$. For the test examples, the normalized prediction from any model was computed as $y^{k}=\frac{\hat{y}^{k}-\mathrm{min}^{k}}{\mathrm{max}^{k}-\mathrm{min}^{k}}$. Then predictions from all models were simply averaged. That is, assuming that there were $K$ models, for any test example, the aggregated clinical prediction was computed as $y^{c}=\frac{1}{K}\sum_{k=1}^{K}y^{k}$.
For pathology models, as the predictions already were in the interval $[0,1]$, we simply averaged predictions from different models. That is, for any test example, the aggregated pathology prediction was computed as $y^{p}=\frac{1}{K}\sum_{1}^{K}y^{p}_{k}$. We generated the final pathology model by averaging the top $K$ models across different combinations of pooling method (mean, max, or MIL), training loss function (Cox or discrete-time), and feature extractor to form the final ensemble. We used $K=10$ for both clinical and pathology models.
@@ 4.4 Digital pathology models
Unlike many prior approaches [43, 32, 44], we did not use any hand-crafted pathomic features. Instead, we obtained learned morphological features from Kestrel, our purpose-built foundation model trained on a pan-cancer dataset. These features were used to train time-to-event models. Below we present our model in greater detail.
@@
Digital pathology slides are extremely high resolution, often exceeding 1 billion pixels per image. Additionally, large regions of these images are empty. Therefore, to extract only meaningful tissue, we used Otsu’s method [45] to distinguish background from foreground. Subsequently, we patchify the foreground into non-overlapping $256\times 256$ patches at 20$\times$ magnification (0.5 microns per pixel). Each slide is thus reduced to a set of patches. There are typically between 8,000 to 16,000 patches per slide.
@@
Slides were preprocessed to extract only foreground patches. These foreground patches were passed through our foundation model to obtain patch-level embeddings. These patch-level embeddings are used as an input to a time-to-event model model.
Our model’s predictions were generated using H&E-stained slides from the same tissue block that was previously used for Oncotype DX testing. For an overwhelming majority of the cases (approximately 93% of the patients across all evaluation cohorts), we used only a single slide. For the remaining minority of the cases for which we had more than one slide, we averaged the predictions across slides.
4.4.1 Training the digital pathology models
We trained two types of models using features extracted by Kestrel as input: Cox proportional hazard models and discrete time models.
@@ 4.4.3 Discrete-time model
For subject $i$ with covariates $x_{i}$, the hazard in interval $A_{j}=(t_{j-1},t_{j}]$ is
@@
and survival probability is then
@@
Then we can model the score at time $j$ as $\phi_{j}(X_{i})$, and to make the hazard between 0 and 1, we set
@@
We can estimate $\phi_{j}(x_{i})$ by minimizing the negative log-likelihood.
Regularization is added to $\mathcal{L_{\text{DT}}}$ the same way it is added to $\mathcal{L_{\text{Cox}}}$ in the previous section.
@@ 4.4.4 Aggregation of patch embeddings
It has also been shown that simple attention-based networks, such as gated attention [47], are also effective ways of aggregating patch embeddings in digital pathology. Thus, we trained a complementary set of time-to-event models using a gated attention network with a discrete-time loss.
@@ 4.5 Clinical models
For building a model for clinical variables we utilized CatBoost [48], a gradient boosting decision tree ensemble. Its primary advantage over other possible models for this application, e.g., neural networks, is how it handles categorical variables and associated missing values which are inevitable in clinical data. Here, CatBoost is specifically trained with the AFT (accelerated failure time) loss. We utilized AFT models with three potential distributions: normal, logistic, and extreme value.
We used eight routinely collected variables (see Table 2) that characterize breast cancer and are typically used to guide treatment decisions. For instance, ER and PR statuses help determine the tumor’s hormone receptor sensitivity, which is critical for hormone-based therapies. HER2 status is important for targeted therapies like trastuzumab, which is effective in HER2+ cancers.
@@ Acknowledgements
JW and KJG contributed to study conception. CFG, LP, YLC and KJG contributed as research advisors. NC, FS, UO, VS, LMP, NK, AB and FJE provided clinical guidance. JW, KZ, JC and KL wrote code, developed infrastructure and trained models throughout the study. JW, KZ, JC, JE, EDC, YJK, FH, ZS, NT, MS, FY, EL, TH, BS, DR, AC, VS, MV, LS, SMS, DB, KC, YZ, LD, AS, WA, RB, JWP, HM, SHT, VA, DZP, AL, BDP, CB, SYJ, JPY, SHL and KJG worked on data preparation. JW, KZ, JC, IO, CFG and KJG performed evaluation and analysis. IO contributed as a statistical advisor. JW, KZ, JC, IO, CFG and KJG worked on drafting and revising the manuscript. All authors critically reviewed the paper and the results and approved the final version.
@@
JW, KZ, JC, CFG, FS, ZS, AB, FJE, YLC and KJG are equity holders of Ataraxis AI. IO served as a consultant for Ataraxis AI. New York University (NYU) maintains financial and intellectual property interests in Ataraxis AI that are pertinent to the research presented in this manuscript. JW and KJG are inventors on a US patent filed corresponding to some of the methodological aspects of this work. The remaining authors declare no competing interests.
@@ A.1 Concordance index
C-index, originally proposed by Harrell et al. [49], is a metric of accuracy for time-to-event models, analogous to AUROC for classification, adjusted for the possibility of censoring. It is defined as the ratio of the number of concordant pairs of test examples to the number of all comparable pairs of test examples. A test example is defined by its event time, $T$, its vector of input features, $x$, and $\delta$ indicating the presence of an event. Two test examples are comparable if the example with the shorter event time is not censored (i.e., experienced an event). Two test examples are concordant if the model assigns a higher risk score to the example with the shorter event time. Given $\delta_{i}=1$ indicating the presence of an event and $T_{i}$ the ‘time-to-event’ response for example $i$, we use Equation 1 to compute the C-index:
@@ A.2 Hazard ratio
Hazard ratio is a metric appropriate for measuring how well a model can divide patients into two groups, e.g. low- and high-risk. The hazard ratio is computed using the Cox proportional hazards model, in which the input indicates one of the two groups. That is, we model the hazard for an individual at time $t$ as $\lambda(t|x)=\lambda_{0}(t)\exp{(\beta x)}$, where $x$ is the group indicator ($x=0$ or $x=1$), $\lambda_{0}(t)$ is the baseline hazard function which represents the hazard when $x=0$ and $\beta$ is the regression coefficient associated with the variable indicating the group a patient belongs to. Then, based on the data, this model is fitted using partial likelihood. For the fitted model, hazard ratio is computed as
@@
If $\mathrm{HR}>1$, the group with $x=1$ has a higher hazard of experiencing the event in comparison to the group with $x=0$. If $\mathrm{HR}<1$, the group with $x=0$ has a higher hazard of experiencing the event in comparison to the group with $x=1$. For example, if $\mathrm{HR}=0.5$ the risk for the group with $x=0$ is twice the risk for the group with $x=1$.
Hazard ratio can be analogously defined for a continuous $x$. Then, it is interpreted as the change in the hazard associated with a one-unit increase in $x$. That is,
@@
Additionally, the logrank test [50] is commonly utilized to determine if there the hazard ratio is significantly different from one between two groups. The Wald test is utilized to assess the significance of individual predictors or the hazard ratio for a specific variable within a multivariate Cox model.
@@ A.3 Random effects model
where $M_{i}$ is the metric for dataset $i$, $\mu$ is the overall effect, $u_{i}$ is the random effect for dataset $i$ ($u_{i}\sim\mathcal{N}(0,\tau^{2})$, $\tau^{2}$ represents the variance between datasets) and $\epsilon_{i}$ is the within-data-set noise ($\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$, $\sigma^{2}$ represents within-data-set variance). The overall fixed effect $\mu$ provides a summary of the model’s average performance across datasets. The model is fit using maximum likelihood, as implemented in the Metagen package [51].
@@ A.4 Endpoint definitions
Table 3: Endpoint definitions and applicable events (local recurrence, distant recurrence, and death from any causes).
@@ A.5 Multivariate Cox analyses
Multivariate analysis was conducted to assess the significance of the pathology model adjusted for the clinical model (see Table 4), importance of AI test after adjusting for datasets (see Table 6), and AI test after adjusting for various additional factors (see Table 6.
Table 4: Cox multivariate analysis was performed across all test sets, evaluating the hazard ratio associated with the clinical and pathology model scores with DFI as the outcome. Separate analyses were conducted for the entire patient cohort and for the subset of patients with Oncotype scores. Dataset indicator variables were included to adjust for differences between datasets.
@@
Table 5: Cox multivariate analysis was conducted across all test sets, evaluating the hazard ratios associated with the AI test with Disease-Free Interval (DFI) as the primary outcome. Dataset-specific indicators were included to adjust for inter-dataset variability in outcomes.
@@
Table 6: Cox multivariate analysis was conducted across all test sets with Oncotype scores, evaluating the hazard ratios associated with the AI test, cancer grade, age, and race, with Disease-Free Interval (DFI) as the primary outcome. Separate analyses were performed with and without the inclusion of the AI score. Dataset-specific indicators were included to adjust for inter-dataset variability in outcomes. To ensure a fair comparison of hazard ratios, we scale both AI test score and Oncotype to a common range of [0-5]. Oncotype, originally on a 0-100 scale, is divided by 20, while our score, on a 0-1 scale, is multiplied by 5.
with AI test without AI test
@@ A.6 Forest plots for all endpoints
Complete results for both primary and exploratory endpoints are available, with C-index shown in Figure 8 and hazard ratios shown in Figure 9. These results include our model’s performance across various endpoints, the performance of the model on individual datasets along with pooled outcomes derived from a random effects model.
@@
Figure 8: Random effect models for aggregating C-index based on continuous AI score for secondary endpoints.
@@
Figure 9: Random effect models for aggregating hazard ratios based on continuous AI score for secondary endpoints.
@@ A.7 Forest plots for clinically relevant subgroups
Complete results different subgroups are available for the primary endpoint of DFI, with C-index shown in Figure 10 and hazard ratios shown in Figure 11. These results include our model’s performance across various endpoints, the performance of the model on individual datasets, along with pooled outcomes derived from a random effects model.
@@
Figure 10: Random effect models for aggregating C-index based on continuous AI score for different patient subgroups.
@@
Figure 11: Random effect models for aggregating hazard ratios based on continuous AI score for different patient subgroups.
@@ A.8 Forest plots for the independent performance of Oncotype DX and the AI test
Full results from the random effects model comparing Oncotype and AI test performance, evaluated in patients with Oncotype scores, are available in Figure 12
@@
Figure 12: Random effect models for comparing the hazard ratios between Oncotype and the AI test for the primary DFI endpoint.

### v2

@@ 4.3.1 Creating model ensembles
We ensembled the clinical models in two steps. First, the output risk scores from all clinical models were scaled approximately to the interval $[0,1]$. That is, for each model $m_{k}$, we found the maximum over the entire validation set, $\mathrm{max}^{k}=\mathrm{max}_{i}\{\hat{y}^{k}_{i}\}$, and the minimum over the entire validation set, $\mathrm{min}^{k}=\mathrm{min}_{i}\{\hat{y}^{k}_{i}\}$. For the test examples, the normalized prediction from any model was computed as $y^{k}=\frac{\hat{y}^{k}-\mathrm{min}^{k}}{\mathrm{max}^{k}-\mathrm{min}^{k}}$. The predictions from all models were then simply averaged. That is, assuming that there were $K$ models, for any test example, the aggregated clinical prediction was computed as $y^{c}=\frac{1}{K}\sum_{k=1}^{K}y^{k}$.
For pathology models, as the predictions were already in the interval $[0,1]$, we simply averaged the predictions from different models. That is, for any test example, the aggregated pathology prediction was computed as $y^{p}=\frac{1}{K}\sum_{1}^{K}y^{p}_{k}$. We generated the final pathology model by averaging the top $K$ models across different combinations of pooling method (mean, max, or multiple instance learning), training loss function (Cox or discrete-time), and feature extractor to form the final ensemble. We used $K=10$ for both clinical and pathology models.
@@ 4.4 Digital pathology models
Unlike many prior approaches [44, 37, 45], we did not use any hand-crafted pathomic features to build the pathology models. Instead, we obtained learned morphological features from Kestrel, our purpose-built foundation model trained on a pan-cancer dataset. These features were used to train time-to-event models. Below we present our model in greater detail.
@@
Digital pathology slides are extremely high resolution, often exceeding 1 billion pixels per image. Additionally, large regions of these images are empty. Therefore, to extract only meaningful tissue, we used Otsu’s method [46] to distinguish background from foreground. Subsequently, we patchify the foreground into non-overlapping $256\times 256$ patches at 20$\times$ magnification (0.5 microns per pixel). Thus, each slide is reduced to a set of patches. There are typically between 8,000 to 16,000 patches per slide.
@@
Slides were preprocessed to extract only foreground patches. These foreground patches were passed through our foundation model to obtain patch-level embeddings. These patch-level embeddings are used as inputs to a time-to-event model.
Our model’s predictions were generated using H&E-stained slides from the same tissue block that was previously used for Oncotype DX testing. For an overwhelming majority of cases (approximately 93% of the patients across all evaluation cohorts), we used only a single slide. For the remaining minority of the cases for which we had more than one slide, we averaged the predictions across slides.
Training the digital pathology models
We trained two types of models using features extracted by Kestrel as input: Cox proportional hazard models and discrete-time models.
@@ 4.4.2 Discrete-time model
For subject $i$ with covariates $x_{i}$ (here, features extracted from a whole slide image), the hazard in interval $A_{j}=(t_{j-1},t_{j}]$ is computed as
@@
and survival probability is computed as
@@
Then, we can model the score in the $j$th time interval using a neural network as $\phi_{j}(x_{i})$. To put the hazard in the $[0,1]$ interval, we set
@@
We can estimate $\phi_{j}(x_{i})$ by minimizing the negative log-likelihood. Regularization is added to $\mathcal{L_{\text{DT}}}$ the same way it is added to $\mathcal{L_{\text{Cox}}}$ described in the previous section.
@@ 4.4.3 Aggregation of patch embeddings
It has been shown that simple attention-based networks, such as gated attention [48], are also effective ways of aggregating patch embeddings in digital pathology. Thus, we trained a complementary set of time-to-event models using a gated attention network with a discrete-time loss.
@@ 4.5 Clinical models
For building a model for clinical variables, we utilized CatBoost [49], a gradient boosting decision tree ensemble. Its primary advantage over other possible models for this application, e.g., neural networks, is how it handles categorical variables which are inevitable in clinical data. Here, CatBoost is specifically trained with the AFT (accelerated failure time) loss [50]. We utilized AFT models with three possible distributions (selected in hyperparameter search): normal, logistic, and extreme value.
We used eight routinely collected variables (see Table 2) that characterize breast cancer and are typically used to guide treatment decisions. For instance, ER and PR statuses help determine tumor hormone receptor sensitivity, which is critical for hormone-based therapies. HER2 status is important for targeted therapies such as trastuzumab, which is effective in HER2+ cancers.
@@ Acknowledgements
JW and KJG contributed to study conception. CFG, LP, YLC and KJG contributed as research advisors. NC, FS, UO, VS, LMP, NK, AB and FJE provided clinical guidance. JW, KZ, JC and KL wrote code, developed infrastructure and trained models throughout the study. JW, KZ, JC, JE, EDC, YJK, FH, ZS, NT, MS, FY, EL, TH, BS, DR, AC, VS, MV, LS, SS, DB, KC, YZ, LD, AS, WA, RB, JWP, HM, SHT, VA, DZP, AL, BDP, CB, SYJ, JPY, SHL and KJG worked on data preparation. JW, KZ, JC, IO, CFG and KJG performed evaluation and analysis. IO contributed as a statistical advisor. JW, KZ, JC, IO, CFG and KJG worked on drafting and revising the manuscript. All authors critically reviewed the paper and the results and approved the final version.
@@
JW, KZ, JC, CFG, FS, ZS, AB, FJE, YLC and KJG are equity holders of Ataraxis AI. IO served as a consultant for Ataraxis AI. New York University (NYU) maintains financial and intellectual property interests in Ataraxis AI that are pertinent to the research presented in this manuscript. JW and KJG are inventors on a US patent application filed corresponding to some of the methodological aspects of this work. The remaining authors declare no competing interests.
@@ A.1 Concordance index
C-index, originally proposed by Harrell et al. [51], is a metric of accuracy for time-to-event models, analogous to AUROC for classification, adjusted for the possibility of censoring. It is defined as the ratio of the number of concordant pairs of test examples to the number of all comparable pairs of test examples. A test example is defined by its event time, $T$, its vector of input features, $x$, and $\delta$ that indicates the presence of an event. Two test examples are comparable if the example with the shorter event time is not censored (i.e., experienced an event). Two test examples are concordant if the model assigns a higher risk score to the example with a shorter event time. Given $\delta_{i}=1$ indicating the presence of an event and $T_{i}$ the ‘time-to-event’ response for example $i$, we use Equation 1 to compute the C-index:
@@ A.2 Hazard ratio
Hazard ratio is a metric appropriate for measuring how well a model can divide patients into two groups, e.g., low- and high-risk. The hazard ratio is computed using the Cox proportional hazards model, in which the input indicates one of the two groups. That is, we model the hazard for an individual at time $t$ as $\lambda(t|x)=\lambda_{0}(t)\exp{(\beta x)}$, where $x$ is the group indicator ($x=0$ or $x=1$), $\lambda_{0}(t)$ is the baseline hazard function which represents the hazard when $x=0$ and $\beta$ is the regression coefficient associated with the variable that indicates the group a patient belongs to. Then, based on the data, this model is fitted using partial likelihood. For the fitted model, hazard ratio is computed as
@@
If $\mathrm{HR}>1$, the group with $x=1$ has a higher hazard of experiencing the event in comparison to the group with $x=0$. If $\mathrm{HR}<1$, the group with $x=0$ has a higher hazard of experiencing the event compared to the group with $x=1$. For example, if $\mathrm{HR}=0.5$ the risk for the group with $x=0$ is twice the risk for the group with $x=1$.
Hazard ratio can be defined analogously for a continuous $x$. Then, it is interpreted as the change in the hazard associated with a one-unit increase in $x$. That is,
@@
Additionally, the logrank test [52] is commonly utilized to determine if the hazard ratio is significantly different from one between two groups. The Wald test is used to assess the significance of individual predictors or the hazard ratio for a specific variable within a multivariate Cox model.
@@ A.3 Random effects model
where $M_{i}$ is the metric for dataset $i$, $\mu$ is the overall effect, $u_{i}$ is the random effect for dataset $i$ ($u_{i}\sim\mathcal{N}(0,\tau^{2})$, $\tau^{2}$ represents the variance between datasets) and $\epsilon_{i}$ is the within-data-set noise ($\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$, $\sigma^{2}$ represents within-dataset variance). The overall fixed effect $\mu$ provides a summary of the model’s average performance across datasets. The model is fit using maximum likelihood, as implemented in the Metagen package [53].
@@ A.4 Endpoint definitions
We selected disease-free interval (DFI) (as defined in Table A.3) as the primary endpoint because it encompasses a broad range of breast cancer-related outcomes. We deemed overall survival and disease-free survival, both of which includes death as part of the endpoint, to be less relevant to our work since not all deaths are related to breast cancer. Ideally, we could exclude all deaths unrelated to breast cancer, but retrospectively determining the cause of death is challenging. Similarly, we deemed distant recurrence to be less informative as it excludes local-regional recurrences, which we also consider to be clinically significant outcomes.
Table A.3: Events (local recurrence, distant recurrence, and/or death from any causes) corresponding to each endpoint.
@@ A.5 Multivariate Cox analyses
Multivariate analysis was conducted to assess the significance of the pathology model adjusted for the clinical model (see Table A.4), importance of AI test after adjusting for datasets (see Table A.5), and AI test after adjusting for various additional factors (see Table A.6).
Table A.4: Cox multivariate analysis was performed across all test sets, evaluating the hazard ratio associated with the clinical and pathology model scores with DFI as the endpoint. Separate analyses were conducted for the entire patient cohort and for the subset of patients with Oncotype scores. Dataset indicator variables were included to adjust for differences between datasets.
@@
Table A.5: Cox multivariate analysis was conducted across all test sets, evaluating the hazard ratios associated with the AI test with Disease-Free Interval (DFI) as the primary endpoint. Dataset-specific indicators were included to adjust for inter-dataset variability in outcomes.
@@
Table A.6: Cox multivariate analysis was conducted across all test sets with Oncotype scores, evaluating the hazard ratios associated with the AI test, cancer grade, age, and race, with Disease-Free Interval (DFI) as the primary endpoint. Separate analyses were performed with and without the inclusion of the AI score. Dataset-specific indicators were included to adjust for inter-dataset variability in outcomes. To ensure a fair comparison of hazard ratios, we scale both AI test score and Oncotype to a common range of [0-5]. Oncotype, originally on a 0-100 scale, is divided by 20, while our score, on a 0-1 scale, is multiplied by 5.
without AI test with AI test
@@ A.6 Forest plots for all endpoints
Here, we present complete results for primary and exploratory endpoints. C-index is shown in Figure A.8 and hazard ratios is shown in Figure A.9. These results include our model’s performance across various endpoints, the performance of the model on individual datasets along with pooled outcomes derived from a random effects model.
@@
Figure A.8: Random effect models for aggregating C-index based on continuous AI score for DFI and secondary endpoints.
@@
Figure A.9: Random effect models for aggregating hazard ratios based on continuous AI score for DFI and secondary endpoints.
@@ A.7 Forest plots for clinically relevant subgroups
Here, we present complete results different subgroups for the primary endpoint of DFI. C-index is shown in Figure A.10 and hazard ratios is shown in Figure A.11. These results include our model’s performance across various endpoints, the performance of the model on individual datasets, along with pooled outcomes derived from a random effects model.
@@
Figure A.10: Random effect models for aggregating C-index based on continuous AI score for different patient subgroups.
@@
Figure A.11: Random effect models for aggregating hazard ratios based on continuous AI score for different patient subgroups.
@@ A.8 Forest plots for the independent performance of Oncotype DX and the AI test
Full results from the random effects model comparing Oncotype and AI test performance, evaluated in patients with Oncotype scores, are presented in Figure A.12
@@
Figure A.12: Random effect models for comparing the hazard ratios between Oncotype and the AI test for the primary DFI endpoint.